The measure of the supplement of and angle is four times the measure of its compliment. Find the measure of the angle.
Answer by stanbon(75887)

The measure of the supplement of and angle is four times the measure of its compliment. Find the measure of the angle.
----------------------
Let the measure of the angle be "x":
EQUATIONs;
180-x = 4(90-x)
180-x = 360-4x
3x = 180
x = 60 degrees
========================
Cheers,
Stan H.